# Who Said Neural Networks Aren’t Linear?

### v1

@@ 3.3 Linear idempotent generative network
Figure 6: Left: IGN training diagram. Black solid arrows denote the forward pass; the red dashed arrow shows backpropagation. The parameter logits are passed through a sigmoid and then thresholded to form the binary projector AA; during backprop, the straight–through estimator (STE) bypasses the threshold so gradients flow through the underlying probabilities. Right: Projection results. Our Linear IGN makes a global projector that projects any input to the target distribution. Top are inputs and bottom are matching outputs.
Idempotency, $f(f(x))=f(x)$, is a central concept in algebra and functional analysis. In machine learning, it has been used in Idempotent Generative Networks (IGNs) (Shocher et al., 2024), to create a projective generative model. It was also used for robust test time training (Durasov et al., 2025). Enforcing a network to be idempotent is tricky and is currently done by using sophisticated optimization methods such as (Jensen & Vicary, 2025). The result is only an approximately idempotent model over the training data. We demonstrate enforcing accurate idempotency through architecture using our Linearizer. The key observation is that idempotency is preserved between the inner matrix $A$ and the full function $f$.
Lemma 8.
@@
$\;\iff\;\big\{\,g^{-1}\circ A\circ\underbrace{(g\circ g^{-1})}_{\mathrm{Id}}\circ A\circ g\,\big\}(x)=\big\{\,g^{-1}\circ A\circ g\,\big\}(x)$
$\;\iff\;f(f(x))=f(x)\quad\text{(by the definition of $f$) \qed}$
@@
Figure 6 left shows the method. Recall that idempotent (projection) matrices have eigenvalues that are either 0 or $1$. We assume a diagonalizable projection matrix $A=Q\Lambda Q^{-1}$ where $\Lambda$ is a diagonal matrix with entries $\{0,1\}$. The matrices $Q,Q^{-1}$ can be absorbed into $g,g^{-1}$ without loss of expressivity. So we can train a Linearizer $g^{-1}(\Lambda g(\cdot))$. In order to have $\Lambda$ that is binary yet differentiable, we use straight-through estimation (Bengio et al., 2013) having underlying probabilities $P$ as parameters in $[0,1]$ and then use A=round(P)+$P$.detach()A=round(P)+P-P.detach() where detachdetach means stopping gradients. We forward propagate rounded values (0 or $1$) and back propagate continuous values.
Idempotency by architecture allows reducing the idempotent loss used in Shocher et al. (2024). We need the data to be fixed points with the tightest possible latent manifold. The losses are:
[…]
$\mathcal{L}_{sparse}$ $=\frac{1}{N}Rank(A)=\frac{1}{N}\sum_{i}^{N}\lambda_{i}$ (22)
We further enforce regularization that encourages $g$ to preserve the norms.
$\mathcal{L}_{isometry}$ $=\Bigl\lVert\lVert g(x)-g(0)\rVert^{2}-\lVert x\rVert^{2}\Bigr\rVert_{1}$ (23)
This nudges $g$ towards a near–isometry around the data, thus resisting mapping far-apart inputs to close outputs. It mitigates collapse and improves stability. We apply it with a small weight ($0.001$).
@@
4 Analysis and Discussion
4.1 Expressiveness and Limitations
A natural question is whether any arbitrary function can be represented by a Linearizer. The answer is no; the architecture imposes certain topological constraints. For example, the null space of $f$ is completely determined by the null space of the matrix $A$. As linear operators can only have null spaces containing a single point (the zero vector) or an infinite subspace, $f$ is restricted to these behaviors. This prevents a Linearizer from, for example, mapping three distinct points to zero.
However, for many practical machine learning tasks, we empirically find that this theoretical limitation is less restrictive than it appears. High-dimensional data, such as images or text embeddings, often reside on a low-dimensional manifold granting the network a large degree of freedom for mapping off-manifold inputs. We hypothesize that in such cases most of the expressiveness constraints may not impact the task for the relevant data. Specifically, regarding the null space example, a finite number of on-manifold inputs can be mapped to zero, while infinitely many more off-manifold.
4.2 On the Role of the Core Operator A
One might also ask whether the matrix $A$ is strictly necessary or if its action could be fully absorbed by the powerful invertible networks $g_{x}$ and $g_{y}$. Let the SVD of $A$ be $U\Sigma V^{T}$. In the general case where $g_{x}$ and $g_{y}$ are distinct, we could define new invertible networks $g^{\prime}_{y}=g_{y}\circ U$ and $g^{\prime}_{x}=V^{T}\circ g_{x}$. The function would then be $f(x)=(g^{\prime}_{y})^{-1}(\Sigma g^{\prime}_{x}(x))$. If we further allow the networks to absorb the scaling defined by the singular values in $\Sigma$, then the core operator could be reduced to a diagonal matrix of zeros and ones (note that this does not imply idempotency unless $g_{x}=g_{y}$). In this sense, the fundamental role of $A$ is simply to define the rank of the function $f$.
The situation is different in the constrained case where $g_{x}=g_{y}=g$. Here, the two networks are not independent. To preserve the structure $f(x)=g^{-1}(Ag(x))$, any transformation absorbed by $g$ must be inverted by $g^{-1}$. This rules out $g$ absorbing SVD elements $U,V$ as they are not necessarily mutual inverses. For diagonalizable matrices however, Eigen-decomposition can be applied. Then, $A=Q\Lambda Q^{-1}$, where one could define $g^{\prime}=Q^{-1}\circ g$ and $f(x)=(g^{\prime})^{-1}(\Lambda g^{\prime}(x))$. In this scenario, $A$ cannot be reduced to a projection matrix, its role is to determine the eigenvalues of the function, a much stronger responsibility than just its rank. However, this is limited to diagonalizable matrices.
In practice, parameterizing an explicit and expressive $A$ is needed in some cases, particularly when we wish to represent a family of functions (e.g., $A_{1},A_{2}\dots$) that share the same linear basis ($g_{x},g_{y}$). The expressiveness of $A$ allows more degrees of freedom for the diversity of different functions. For example, in our diffusion application, we learn a time-dependent operator $A_{t}$. A single pair of networks, $g_{x}$ and $g_{y}$, defines a fixed basis, while explicit $A_{t}$ allows for a rich variety of linear operations within that basis, controlled by the continuous parameter $t$.
@@ 6 Related Work
Linearization in Dynamical Systems: The Koopman operator theory linearizes nonlinear dynamics by advancing observables on a single state space, yielding linear evolution $z_{t+1}=Kz_{t}$ in discrete time or $\dot{z}=Az$ in continuous time (Mezić, 2005). Data-driven realizations such as DMD and EDMD construct finite-dimensional approximations of this infinite-dimensional operator on a chosen dictionary of observables, which typically produces a square matrix because the dictionary is mapped into itself, although variants with different feature sets exist (Schmid, 2010; Lusch et al., 2018; Mardt et al., 2018). Linear algebraic analyses such as SVD, eigendecomposition, and taking powers are standard in these Koopman approximations. Our work addresses a different object: an arbitrary learned mapping $f:\mathcal{X}\to\mathcal{Y}$. We learn invertible coordinate maps $g_{x}$ and $g_{y}$ and a finite matrix $A$ such that $g_{y}\circ f\circ g_{x}^{-1}=A$, which gives the exact finite-dimensional linearity of $f$ between induced spaces. This formulation allows for distinct input and output coordinates, allows $A$ to be rectangular when $\dim\mathcal{X}\neq\dim\mathcal{Y}$, and transfers linear-algebraic structure in $A$ directly to $f$ in a controlled way. In particular, one can compute $\mathrm{SVD}(f)$ via $\mathrm{SVD}(A)$, perform analytic test-time inversion with $A^{+}$, impose spectral constraints such as idempotency by restricting $\sigma(A)$, and more.
@@
Neural Tangent Kernel: The NTK framework shows that infinitely wide networks trained with small steps evolve linearly in parameter space (Jacot et al., 2018). However, the resulting model remains nonlinear in input, output mapping. In comparison, our contribution achieves input-output linearity in a learned basis for any network width, independent of training dynamics.
@@
One-Step Diffusion Distillation: Reducing sampling cost in diffusion models often uses student distillation: Progressive Distillation (Salimans & Ho, 2022), Consistency Models (Song et al., 2023), Distribution Matching Distillation (DMD) and f-distillation fall into this category. A recent work, the Koopman Distillation Model (KDM) (Berman et al., 2025), uses Koopman-based encoding to distill a diffusion model into a one-step generator, achieving strong FID improvements through offline distillation. The Koopman Distillation Model (KDM) distills a pre-trained diffusion teacher into an encoder linear decoder for diffusion only. We train from scratch, enforce exact finite-dimensional linearity by construction, and target multiple use cases including composition, inversion, idempotency, and continuous evolution. Unlike some invertible Koopman setups that force a shared map, KDM employs distinct encoder components and a decoder and does not enforce $g_{x}=g_{y}$.
@@
Invertible Neural Networks: Invertible models like NICE (Dinh et al., 2015), RealNVP (Dinh et al., 2017) and Glow (Kingma & Dhariwal, 2018) use invertible transforms for density modeling and generative sampling. In this work, we make use of such invertible neural networks, to impose a bijective change of coordinates that allows linearity (lemma 2 only holds given this invertability of $g_{x},$ $g_{y}$) rather than modeling distribution.
@@
Group-theoretic and Equivariant Representations: Let $G$ act on $\mathcal{X}$ via $(g,x)\mapsto g\cdot x$. Many works learn an encoder $e:\mathcal{X}\to V$ and a linear representation $\rho:G\to\mathrm{GL}(V)$ such that the equivariance constraint $e(g\cdot x)=\rho(g)\,e(x)$ holds approximately, often with a decoder $d$ that reconstructs $x$ or enforces $d(\rho(g)z)\approx g\cdot d(z)$ (Quessard et al., 2020). The objective is to make the action of a symmetry group linear in latent coordinates. In contrast, we target a different object: an arbitrary learned map $f:\mathcal{X}\to\mathcal{Y}$. We construct invertible $g_{x},g_{y}$ and a finite matrix $A$ so that $g_{y}\circ f\circ g_{x}^{-1}=A,$ which yields exact finite-dimensional linearity of $f$ between induced spaces and enables direct linear algebra on $f$ via $A$.
@@
Manifold Flattening and Diffeomorphic Autoencoders: Manifold-flattening methods assume data lie on a manifold $M\subset\mathbb{R}^{N}$ and learn or construct a map $\Phi:\mathbb{R}^{N}\to\mathbb{R}^{k}$ so that $\Phi(M)$ is close to a linear subspace $L$, often with approximate isometry on $M$ (Psenka et al., 2024). Diffeomorphic autoencoders parameterize deformations $\varphi\in\mathrm{Diff}(\Omega)$ with a latent $z$ in a Lie algebra and use a decoder to warp a template, with variants using a log map to linearize the deformation composition (Bône et al., 2019). These approaches flatten data geometry or deformation fields. This contrasts with our Linearizer in which the flattening is over a mapping function rather than a dataset manifold or a deformation group.
@@
Deep Linear Networks: Networks composed solely of linear layers are linear in the Euclidean basis and are primarily used to study optimization paths (Saxe et al., 2014; Arora et al., 2018; Cohen et al., 2016). They lack expressive power in standard coordinates. In contrast, our Linearizer is expressive thanks to $g_{x}$ and $g_{y}$ that are nonlinear over the standard vector spaces, yet maintains exact linearity in the induced coordinate system.
@@
Normalizing flows: Normalizing flows Rezende & Mohamed (2015) use invertible neural networks to transport probability measures: they map a simple Gaussian distribution into a complex data distribution while preserving density through the change-of-variables formula. Linearizers use the same building block, an invertible network, but instead of transporting measures, they transport algebraic structure. In this sense, Linearizers may be viewed as the “linear-algebraic analogue” of normalizing flows: both frameworks exploit invertible transformations to pull back complex objects into domains where they become simple and tractable.
@@ 7 Conclusion
We introduced the Linearizer, a framework that learns invertible coordinate transformations such that neural networks become exact linear operators in the induced space. This yields new applications in one-step flow matching, modular style transfer, and architectural enforcement of idempotency.
Looking ahead, several directions stand out. First, scaling one-step flow matching and IGN to larger datasets and higher resolutions could provide competitive generative models with unprecedented efficiency. Second, the matrix structure of the Linearizer opens the door to modeling motion dynamics: by exploiting matrix exponentials, one could simulate continuous-time evolution directly, extending the approach beyond generation to physical and temporal modeling. More broadly, characterizing the theoretical limits of Linearizer expressivity, and integrating it with other operator-learning frameworks, remains an exciting avenue for future research.
Acknowledgements
We thank Amil Dravid and Yoad Tewel for insightful discussions and for sharing ideas that helped shape this work. A.S. is supported by the Chaya Career Advancement Chair.
Appendix A Proof of Lemma 1 (Valid vector spaces)
@@
Appendix B Proof of lemma 6 (SVD of a Linearizer)
@@
Appendix C Proof of Lemma 4 (Hilbert Space)
@@
By Lemma 1, $(V,\oplus_{g},\odot_{g})$ is a vector space and $g$ is a vector-space isomorphism from $(V,\oplus_{g},\odot_{g})$ onto $(\mathbb{R}^{n},+,\cdot)$. The form $\langle\cdot,\cdot\rangle_{g}$ is an inner product because it is the pullback of the Euclidean inner product: bilinearity, symmetry, and positive definiteness follow immediately from injectivity of $g$ and the corresponding properties in $\mathbb{R}^{n}$.
@@
Appendix D Proof of Lemma 7 (Pseudo-Inverse of a Linearizer)

### v2

@@ 4.3 Linear Idempotent Generative Network
Figure 6: Left: IGN training diagram. Black solid arrows denote the forward pass; the red dashed arrow shows backpropagation. The parameter logits are passed through a sigmoid and then thresholded to form the binary projector AA; during backprop, the straight-through estimator (STE) bypasses the threshold so gradients flow through the underlying probabilities. Right: Projection results. Our Linear IGN makes a global projector that projects any input to the target distribution. Top are inputs and bottom are matching outputs.
Idempotency, $f(f(x))=f(x)$, is a central concept in algebra and functional analysis. In machine learning, it has been used in Idempotent Generative Networks (IGNs) (Shocher et al., 2024), to create a projective generative model. It was also used for robust test time training (Durasov et al., 2025). Enforcing a network to be idempotent is tricky and is currently done by using sophisticated optimization methods such as (Jensen and Vicary, 2025). The result is only an approximately idempotent model over the training data. We demonstrate enforcing accurate idempotency through architecture using our Linearizer. The key observation is that idempotency is preserved between the inner matrix $A$ and the full function $f$.
Lemma 4.2.
@@
$\;\iff\;\big\{\,g^{-1}\circ A\circ\underbrace{(g\circ g^{-1})}_{\mathrm{Id}}\circ A\circ g\,\big\}(x)$
$\quad\;=\big\{\,g^{-1}\circ A\circ g\,\big\}(x)$
$\;\iff\;f(f(x))=f(x)$
$\quad\;\text{(by the definition of $f$)}\qed$
@@
Figure 6 left shows the method. Recall that idempotent (projection) matrices have eigenvalues that are either 0 or $1$. We assume a diagonalizable projection matrix $A=Q\Lambda Q^{-1}$ where $\Lambda$ is a diagonal matrix with entries $\{0,1\}$. The matrices $Q,Q^{-1}$ can be absorbed into $g,g^{-1}$ without loss of expressivity. So we can train a Linearizer $g^{-1}(\Lambda g(\cdot))$. In order to have $\Lambda$ that is binary yet differentiable, we use an estimation (Bengio et al., 2013) having underlying probabilities $P$ as parameters in $[0,1]$ and then use $A=\text{round}(P)+P-P.\text{detach}()$ where detach means stopping gradients. We forward propagate rounded values (0 or $1$) and back propagate continuous values.
Idempotency by architecture allows reducing the idempotent loss used in (Shocher et al., 2024). We need the data to be fixed points with the tightest possible latent manifold. The losses are:
$\mathcal{L}_{\text{rec}}$ $=\|f(x)-x\|^{2}$ (20)
$\mathcal{L}_{\text{sparse}}$ $=\frac{1}{N}\text{Rank}(A)=\frac{1}{N}\sum_{i}^{N}\lambda_{i}$ (21)
We further enforce regularization that encourages $g$ to preserve the norms:
$\mathcal{L}_{\text{isometry}}$ $=\Bigl\lVert\lVert g(x)-g(0)\rVert^{2}-\lVert x\rVert^{2}\Bigr\rVert_{1}$ (22)
This nudges $g$ towards a near-isometry around the data, thus resisting mapping far-apart inputs to close outputs. It mitigates collapse and improves stability. We apply it with a small weight ($0.001$).
@@
4.4 Additional Experiments
We provide more experiments of diverse types of data and tasks to empirically test the robustness of the Linearizer and its ability to learn various mappings.
CelebFaces Attributes (CelebA).
We evaluate on the CelebA face-attribute classification task, which includes large pose variation, background clutter, diverse identities, and rich annotations (40 binary attributes such as Male, Eyeglasses, Blond Hair). After training, our Linearizer attains a test accuracy of 87.6%; for reference, a ResNet-50 baseline reaches 86.1%. These results suggest that the Linearizer handles moderately complex, real-world visual variability.
Distillation of nonlinear disentanglement models.
In this experiment, we distill the encoder of a pretrained disentanglement model into a Linearizer (using pretrained checkpoints). We consider VAE, FactorVAE, and $\beta$-VAE on the dSprites dataset, which provides ground-truth factors of variation. We assess quality with standard metrics: FactorVAE, SAP, and DCI (disentanglement, completeness, informativeness). Across models, the Linearizer closely matches the original encoders (rows prefixed with Linearizer-) on these metrics (often slightly improving FactorVAE/SAP, e.g., BetaH: +0.10/+0.01), indicating that it can capture highly nonlinear transformations with no loss in disentanglement quality of the encoded representations. For the table with full results please see Appendix E.3.
Weather Prediction.
To further explore the applicability of our framework, we conduct a small experiment in a new domain: regression of future steps. We use a standard weather-prediction benchmark where the model must predict the future given 96 past time steps. We compare our method with two well-known baselines tailored for regression, Autoformer (Wu et al., 2021) and PatchTST (Nie, 2022). Briefly, although our method is generic, it shows competitive results, surpassing the highly nonlinear Autoformer and being comparative (e.g., gaps of $\sim$0.008–0.034 MSE depending on horizon) with the strong PatchTST baseline. For the table with full results please see Appendix E.3.
@@ 6 Related Work
Koopman theory linearizes nonlinear dynamics via observables, yielding $z_{t+1}=Kz_{t}$ (discrete) or $\dot{z}=Az$ (continuous) (Mezić, 2005). Data-driven realizations such as DMD and EDMD construct finite-dimensional approximations of this infinite-dimensional operator on a chosen dictionary of observables, which typically produces a square matrix because the dictionary is mapped into itself, although variants with different feature sets exist (Schmid, 2010; Lusch et al., 2018; Mardt et al., 2018). Linear algebraic analyses such as SVD, eigendecomposition, and taking powers are standard in these Koopman approximations. Our work addresses a different object: an arbitrary learned mapping $f:\mathcal{X}\to\mathcal{Y}$. We learn invertible coordinate maps $g_{x}$ and $g_{y}$ and a finite matrix $A$ such that $g_{y}\circ f\circ g_{x}^{-1}=A$, which gives the exact finite-dimensional linearity of $f$ between induced spaces. This formulation allows for distinct input and output coordinates, allows $A$ to be rectangular when $\dim\mathcal{X}\neq\dim\mathcal{Y}$, and transfers linear-algebraic structure in $A$ directly to $f$ in a controlled way. One can compute $\mathrm{SVD}(f)$ via $\mathrm{SVD}(A)$, invert analytically with $A^{+}$, impose spectral constraints (e.g., idempotency) via $\sigma(A)$, and more.
Neural Tangent Kernel.
@@
The NTK framework shows that infinitely wide networks trained with small steps evolve linearly in parameter space (Jacot et al., 2018). However, the resulting model remains nonlinear in input-output mapping. In comparison, our contribution achieves input-output linearity in a learned basis for any network width, independent of training dynamics.
One-Step Diffusion Distillation.
@@
Reducing sampling cost in diffusion models often uses student distillation: Progressive Distillation (Salimans and Ho, 2022), Consistency Models (Song et al., 2023), Distribution Matching Distillation (DMD) and f-distillation fall into this category. A recent work, the Koopman Distillation Model (KDM) (Berman et al., 2025), uses Koopman-based encoding to distill a diffusion model into a one-step generator, achieving strong FID improvements through offline distillation. The Koopman Distillation Model (KDM) distills a pre-trained diffusion teacher into an encoder-linear-decoder for diffusion only. We train from scratch, enforce exact finite-dimensional linearity by construction, and target multiple use cases including composition, inversion, idempotency, and continuous evolution. Unlike some invertible Koopman setups that force a shared map, KDM employs distinct encoder components and a decoder and does not enforce $g_{x}=g_{y}$.
Invertible Neural Networks.
@@
Invertible models like NICE (Dinh et al., 2015), RealNVP (Dinh et al., 2017) and Glow (Kingma and Dhariwal, 2018) use invertible transforms for density modeling and generative sampling. In this work, we make use of such invertible neural networks to impose a bijective change of coordinates that allows linearity (Proposition 2.4 only holds given this invertibility of $g_{x}$, $g_{y}$) rather than modeling distribution.
Group-theoretic and Equivariant Representations.
@@
Let $G$ act on $\mathcal{X}$ via $(g,x)\mapsto g\cdot x$. Many works learn an encoder $e:\mathcal{X}\to V$ and a linear representation $\rho:G\to\mathrm{GL}(V)$ such that the equivariance constraint $e(g\cdot x)=\rho(g)\,e(x)$ holds approximately, often with a decoder $d$ that reconstructs $x$ or enforces $d(\rho(g)z)\approx g\cdot d(z)$ (Quessard et al., 2020). The objective is to make the action of a symmetry group linear in latent coordinates. In contrast, we target a different object: an arbitrary learned map $f:\mathcal{X}\to\mathcal{Y}$. We construct invertible $g_{x},g_{y}$ and a finite matrix $A$ so that $g_{y}\circ f\circ g_{x}^{-1}=A$, which yields exact finite-dimensional linearity of $f$ between induced spaces and enables direct linear algebra on $f$ via $A$.
Manifold Flattening and Diffeomorphic Autoencoders.
@@
Manifold-flattening methods assume data lie on a manifold $M\subset\mathbb{R}^{N}$ and learn or construct a map $\Phi:\mathbb{R}^{N}\to\mathbb{R}^{k}$ so that $\Phi(M)$ is close to a linear subspace $L$, often with approximate isometry on $M$ (Psenka et al., 2024). Diffeomorphic autoencoders parameterize deformations $\varphi\in\mathrm{Diff}(\Omega)$ with a latent $z$ in a Lie algebra and use a decoder to warp a template, with variants using a log map to linearize the deformation composition (Bône et al., 2019). These approaches flatten data geometry or deformation fields, whereas our Linearizer flattens a mapping function rather than a manifold or deformation group.
Deep Linear Networks.
@@
Networks composed solely of linear layers are linear in the Euclidean basis and are primarily used to study optimization paths (Saxe et al., 2014; Arora et al., 2018; Cohen et al., 2016). They lack expressive power in standard coordinates. In contrast, our Linearizer is expressive thanks to $g_{x}$ and $g_{y}$ that are nonlinear over the standard vector spaces, yet maintains exact linearity in the induced coordinate system.
Normalizing Flows.
@@
Normalizing flows (Rezende and Mohamed, 2015) use invertible networks to map Gaussians to complex data distributions, preserving density via the change-of-variables formula. Linearizers use the same building block, an invertible network, but instead of transporting measures, they transport algebraic structure. In this sense, Linearizers may be viewed as the “linear-algebraic analogue” of normalizing flows: both frameworks exploit invertible transformations to pull back complex objects into domains where they become simple and tractable.
@@ 7 Conclusion
We introduced the Linearizer, a framework that learns invertible coordinate transformations such that neural networks become exact linear operators in the induced space. This yields new applications in one-step flow matching, modular style transfer, and architectural enforcement of idempotency. Looking ahead, several directions stand out. First, scaling one-step flow matching and IGN to larger datasets and higher resolutions could provide competitive generative models with unprecedented efficiency. Second, the matrix structure of the Linearizer opens the door to modeling motion dynamics: by exploiting matrix exponentials, one could simulate continuous-time evolution directly, extending the approach beyond generation to physical and temporal modeling. More broadly, characterizing the theoretical limits of Linearizer expressivity, and integrating it with other operator-learning frameworks, remains an exciting avenue for future research.
Impact Statement
This paper presents a fundamental architectural framework intended to advance the theoretical understanding and computational efficiency of neural networks. A primary application of our work is accelerating generative diffusion models via one-step sampling. While this offers significant benefits in terms of reduced computational cost and energy consumption during inference, it shares the general societal risks associated with generative AI, such as the potential for creating misleading synthetic content. We believe our specific method does not introduce new ethical concerns beyond those already established in the field of generative modeling.
Appendix A Proof of Proposition 2.3 (Valid Vector Spaces)
@@
Appendix B Proof of Proposition 2.9 (SVD of a Linearizer)
@@
Appendix C Proof of Proposition 2.7 (Hilbert Space)
@@
By Proposition 2.3, $(V,\oplus_{g},\odot_{g})$ is a vector space and $g$ is a vector-space isomorphism from $(V,\oplus_{g},\odot_{g})$ onto $(\mathbb{R}^{n},+,\cdot)$. The form $\langle\cdot,\cdot\rangle_{g}$ is an inner product because it is the pullback of the Euclidean inner product: bilinearity, symmetry, and positive definiteness follow immediately from injectivity of $g$ and the corresponding properties in $\mathbb{R}^{n}$.
@@
Appendix D Proof of Lemma 4.1 (Pseudo-Inverse of a Linearizer)